OpenStudy (anonymous):

How do you simplify √343

OpenStudy (anonymous):

to simplify a radical,write the radicand as a product of a perfect square and another number,ex \[\sqrt{32}=\sqrt{16x2}=\sqrt{16}x \sqrt{2}=4\sqrt{2}\]

OpenStudy (anonymous):

thank you! To check my work is the answer 7√7?